PART A: If you invest $11,000 today at 11% interest, how much will you have in...

PART A:

If you invest $11,000 today at 11% interest, how much will you have in 11 years?

A. $28,016

B. $17,688

C. $34,669

D. $31,712

PART B:

How much must you invest at 5% interest in order to see your investment grow to $5,000 in 5 years?

A. $3,918
B. $3,415
C. $3,105
D. None of the above
Answer #1

Solution :

Part A )

Principal = $11,000

Interest rate = 11%

Time = 11 year

We will use compound interest formula to calculate the future value

FV = Principal * ( 1+ interest rate ) ^ years

FV = 11,000 * ( 1+11% ) ^11 = 11,000* 3.151757 = $34,669

Correct option is C )

Part B )

Here it is given that interest rate = 5% , FV = 5000, year = 5 . We need to find the principal value

5000 = principal * ( 1+0.05) ^5

Principal = 5000/ 1.276282 = 3917.63 = 3918

Correct option is A )